Find the pressure increase in the fluid in a syringe when a nurse applies a force of 45N to the syringe circular piston, which has a radius of 1.2cm.
1
Expert's answer
2020-07-21T13:48:44-0400
"P=\frac{F}{A}=\frac{F}{\pi r^2}\\P=\frac{45}{\pi (0.012)^2}=99500\ Pa"